OpenStudy (precal):

Find the value f ' (0) to determine if f(x) is differentiable at x=0

OpenStudy (precal):

\[f(x)=x ^{1/3}+2\]

OpenStudy (precal):

ok first I was asked to graph the function second I was asked if f(x) is continuous at x=0

OpenStudy (precal):

so I did the graph, no problem yes f(x) is continuous at x=0

OpenStudy (precal):

I took the derivative and the derivative should be undefined

myininaya (myininaya):

is f continuous at 0? or is f right continuous at 0?

OpenStudy (precal):

what do you mean f right continuous at o?

myininaya (myininaya):

isn't actually left continuous nothing exists to the left for it to be continuous at x=a you must have: 1)\[\lim_{x \rightarrow a}f(x) \text{ exists }\] 2) \[f(a) \text{ exists }\] finally 3) \[\lim_{x \rightarrow a} f(x)=f(a)\] can you answer number 1?

myininaya (myininaya):

oh is that not a 1/2 power?

myininaya (myininaya):

I cannot read latex I swear

myininaya (myininaya):

3 and 2 look like

OpenStudy (anonymous):



OpenStudy (anonymous):

the power is (1/3-1)

myininaya (myininaya):

but also song that does not give f' is 0 at x=0 \[\frac{1}{3x^\frac{2}{3}}\]

OpenStudy (precal):

I thought I was looking at the wrong graph

OpenStudy (precal):

it is x to the (1/3) power plus 2

myininaya (myininaya):

Yea sorry @precal I thought I was seeing 1/2 not 1/3 :( my bad

myininaya (myininaya):

i have to take my cat to the vet i will bb

OpenStudy (precal):

ok here is where I got very confused. I did this by hand and then I used my graphing calculator to verify the solution and it gave me 100 well that is very confusing or maybe I found a function where the calculator gives an incorrect solution

OpenStudy (precal):

thanks @myininaya hope your cat feels better. Mine never like going to the vet. She did live for about 20 years. The best cat ever.....

OpenStudy (anonymous):

the first thing is need to prove (right =left ) continuous ,and then diferenationate

OpenStudy (precal):

well the f(0)=2 the point exists, and the limit exists at 2 approaching zero from both sides

OpenStudy (anonymous):

yes

OpenStudy (precal):

no f ' (0) does not equal zero that implies that f ' (0) has a horizontal tangent at x=0 and it clearly does not

OpenStudy (precal):

this is the graph of the derivative